A triangle has sides in ratio of 6:8:12. If the the perimeter is 52, what are the dimensions of the triangle?
Answer by jim_thompson5910(35256)

6x + 8x + 12x = 52

26x = 52

x = 52/26

x = 2

-------------------------------------------------------

6x = 6*2 = 12

8x = 8*2 = 16

12x = 12*2 = 24

The dimensions of the triangle are: 12, 16, 24